This is on absolute value...
-4|x-2|-9=-37
I have tried this and I got 9, -5. All I want to do is double check to see if these two answers are correct.

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52831)

.
-4|x-2| - 9 = -37  ====>  -4|x-2| = -37 + 9 = - 28  ====>  


    divide by (-4) both sides. You will get an equivalent equation


|x-2| = 7.


The solutions are TWO numbers that are remoted on the number line in 7 units from 2.

These numbers are  2+7 = 9  and  2-7 = -5.


Answer.  The given equation has two solutions: x= 9  and  x= -2.

YOUR ANSWER IS CORRECT !   Congratulations !!


Answer by Alan3354(69443)

-4|x-2|-9=-37
-4|x-2| = -28
|x-2| = 7
---------
x-2 = 7
x = 9
=================
-(x-2) = 7
x-2 = -7
x = -5